OpenStudy (anonymous):

prove the following trigonometric identities. Include reason with calculation: secθ/tanθ=cscθ

OpenStudy (anonymous):

sec θ / tan θ = csc θ sec θ ÷ tan θ = . 1. . . . . . sin θ --------- ÷ ---------- = cos θ . . . cos θ . 1. . . . . . cos θ . . . . 1 --------- X ---------- = -------- = csc θ cos θ . . . sin θ . . . sin θ

OpenStudy (anonymous):

\[\frac{\sec(\theta)}{\tan(\theta)}=\frac{1}{\cos(\theta)} \times \frac{\cos(\theta)}{\sin(\theta)}=\frac{1}{\sin(\theta)}=\csc(\theta)\]